OpenStudy (yumeybitaa):

Frank needs to fill juice cups for his little sister's birthday party. There are 35 juice cups and each juice cup holds 12 fluid ounces. The juice comes in 1-gallon bottles. How many 1-gallon bottles of juice will Frank need to purchase? (1 quart = 32 fluid ounces, 1 gallon = 4 quarts) 1 bottle 2 bottles 3 bottles 4 bottles

OpenStudy (anonymous):

1 gallon = 4 quarts 1 quart = 32 ounces hence meaning that 1 gallon = 4*32 ounces = 128 ounces so 1 gallon can hold 128 fluid ounces of juice there are 35 juice cups needed, each holding 12 fluid ounces, so the total fluid ounces is 35*12 = 420 fluid ounces so we need 420 fluid ounces, and 1 gallon can give us 128 ounces, so how many gallons would we need to make 420? ---> 420/128 :)